Is the answer for this problem SQRT(5)\sqrt(11) = sqrt(55)\11
Found 2 solutions by CharlesG2, Alan3354:
Answer by CharlesG2(834)

Is the answer for this problem SQRT(5)\sqrt(11) = sqrt(55)\11 ?
sqrt(5)/sqrt(11) (times a factor of 1)
sqrt(5)/sqrt(11) * sqrt(11)/sqrt(11)
(sqrt(5)*sqrt(11))/11
sqrt(55)/11
yes, sqrt(5)/sqrt(11) = sqrt(55)/11

Answer by Alan3354(69443)

sqrt(5)/sqrt(11) does = sqrt(55)/11